OpenStudy (anonymous):

Curtains are manufactured in 3 different styles, 5 different colors, and 4 different fabrics. How many different style-color-fabric combinations are possible?

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

the counting principle says that if there are \(m\) ways to do one thing and \(n\) ways to do another, then there are \(m\times n\) ways to do them together your job is to multiply \(3\times 5\times 4\)

OpenStudy (rina.r):

yeah multiply